What’s the percentage of pbcl2 ?
1
Expert's answer
2019-04-08T07:55:00-0400

Solution.

PbCl2

M(PbCl2) = 278,11 g/mole

Firstly, we find the mass fraction of lead by the following formula:

w(Pb) = (1*207,2*100%)/278,11 = 74,50%

w(Cl) = 100%-74,50% = 25,50%

Answer:

w(Pb) = (1*207,2*100%)/278,11 = 74,50%

w(Cl) = 100%-74,50% = 25,50%